OpenStudy (anonymous):

each of my daughters has as many brothers as sisters, but each one of my sons has three times as many sisters as brothers, how many children do i have?

OpenStudy (kymber):

Mind blown

OpenStudy (inkyvoyd):

Ok

OpenStudy (inkyvoyd):

x is the number of daughters y is the number of sons

OpenStudy (inkyvoyd):

x+1=y

OpenStudy (inkyvoyd):

3y=x

OpenStudy (inkyvoyd):

and, y=-2, so I did something wrong.

OpenStudy (inkyvoyd):

*-1/2

OpenStudy (inkyvoyd):

x+1=y 3(y-1)=x

OpenStudy (anonymous):

\[nd-1=ns\] ns-1=3*nd

OpenStudy (inkyvoyd):

I got it wrong again.

OpenStudy (anonymous):

ns is the number of sons

OpenStudy (anonymous):

\[3*(ns-1)=nd\] instead ns-1=3*nd

OpenStudy (anonymous):

ns=2 and nd=3

OpenStudy (anonymous):

 lol